# Higher-order homogenised riblet boundary conditions

### v1

@@ Abstract
The description of riblets and other drag-reducing devices has long used the concept of longitudinal and transverse protrusion heights, both as a means to predict the drag reduction itself and as equivalent boundary conditions to simplify numerical simulations by transferring the effect of riblets onto a flat virtual boundary. The limitation of this idea is that it stems from a first-order approximation in the riblet-size parameter $s^{+}$, and as a consequence it cannot predict other than a linear dependence of drag reduction upon $s^{+}$; in other words, the initial slope of the drag-reduction curve. Here the concept is extended to a full asymptotic expansion using matched asymptotics, which consistently provides higher-order protrusion heights and higher-order equivalent boundary conditions on a virtual flat surface. This procedure also allows us to explore nonlinearities of the Navier-Stokes equations and the way they enter the $s^{+}$-expansion, with somewhat surprising results.
@@ 1 Introduction
Surfaces that are not smooth alter the flow-induced transfer of momentum, heat and mass (Jiménez, 2004; Bottaro, 2019; García-Mayoral et al., 2019; Chung et al., 2021). Examples include riblets (García-Mayoral & Jiménez, 2011a), shark denticles (Lang et al., 2008; Savino & Wu, 2024), seal fur (Itoh et al., 2006; de Segura & García-Mayoral, 2019), superhydrophobic or liquid-infused surfaces (García-Cartagena et al., 2019), roughness (Busse et al., 2017; Thakkar et al., 2018; Flack & Schultz, 2010, 2014) and wavy surfaces (Charru et al., 2013; Thorsness et al., 1978). Understanding the relationship between form (shape, topography, morphology) and function (alteration of transfer properties) of these surfaces is not trivial, generally requiring challenging laboratory experiments (Abrams & Hanratty, 1985; Bechert et al., 1997; Schultz & Flack, 2013) or numerical simulations (Busse et al., 2017; García-Mayoral & Jiménez, 2011b). However, in regimes where a small parameter exists, inroads can be made with asymptotic analyses.
@@
As discussed in Luchini (2013), surface texture characterized by a small scale can be analysed in two separate distinguished limits, one of which is a shallow-roughness limit where height of the texture tends to zero while its horizontal scales stay fixed. These are shallow or wavy surfaces (Luchini & Charru, 2019; Abrams & Hanratty, 1985). The other limit is the small-roughness limit where all three dimensions tend to zero proportionally to each other, i.e. the texture stays geometrically self-similar, e.g. riblets (García-Mayoral & Jiménez, 2011b; Endrikat et al., 2021; Modesti et al., 2021) and rough surfaces (Thakkar et al., 2018; Sharma & García-Mayoral, 2020) with small surface features compared to flow features. In this paper we shall be concerned with the small-roughness limit. In particular, we will develop a higher-order matched asymptotic expansion that encapsulates the relationship between flow alteration and surface details via equivalent (homogenised) boundary conditions. Apart from their direct use for informing design (e.g. García-Mayoral & Jiménez, 2011a), such boundary conditions circumvent the need to resolve the small surface features in flow simulations, easing a bottleneck that enables routine computations and subsequent analysis towards overall drag savings, for example (Lācis et al., 2020; Ahmed & Bottaro, 2025).
@@
In the context of riblets, Bechert & Bartenwerfer (1989) first developed the idea of the longitudinal protrusion height, which can also be interpreted as a slip length (cf. Bottaro, 2019). When applied in the turbulent-flow regime, it predicts the wall-normal distance below a reference plane (often taken as the plane of the riblet tips) where the mean (longitudinal) flow extrapolates to zero. In this sense, the theory predicts that the mean flow is shifted by riblets, an idea which has been confirmed in numerical simulations (García-Mayoral & Jiménez, 2011b). In the presence of three-dimensional turbulence, it was also necessary to consider the other (non-longitudinal) flow components. This was explored by Luchini et al. (1991), who identified the shift in turbulence to be the lateral- or cross-flow (spanwise to the mean) protrusion height. Then, only can the drag change that modifies the inner layer of wall turbulence be explained and extrapolated to full scale (Luchini, 1996). Specifically, the drag change is proportional to the difference between the longitudinal and cross-flow protrusion heights, which is independent of the reference plane from which the protrusion heights are measured.
The protrusion heights describe well the asymptotic change in drag for vanishingly small roughness in turbulence flow (Wong et al., 2024), however, the effect of wall transpiration, i.e. wall-normal component of motion, needs to also be accounted for where small roughness features are no longer vanishingly small (Ibrahim et al., 2021; Bottaro, 2019). For riblets that are comparable in size to the quasi-streamwise vortices of turbulence, a physically based so-called viscous vortex model was developed by Wong et al. (2024) to allow for this wall-normal component. Another approach is the transpiration-resistance model of Lācis et al. (2020); Khorasani et al. (2022), which are essentially homogenised boundary conditions for the macroscopic flow, with coefficients obtained from microscopic flow calculations performed in a single repeating unit of surface texture. The success of this approach relies on the linearity between the macrosopic flow and the microscopic response, as allowed by Stokes flow. A similar approach is also pursued by Zampogna et al. (2019). To extend to inertial effects within the texture, Zampogna & Bottaro (2016) added an Oseen-like approximation of the advection term, which retains the linearity between the macroscopic forcing the microscopic response. Ahmed & Bottaro (2025) demonstrated the utility of these boundary conditions in direct numerical simulations of turbulence.
@@
A formal higher-order matched-asymptotic expansion of the wall-boundary condition of a periodic surface texture has not been attempted. Although the effect of wall transpiration has been recognised as a second-order effect and modelled to varying degrees (Lācis et al., 2020; Ibrahim et al., 2021; Wong et al., 2024; Ahmed & Bottaro, 2025), it is unclear whether other effects that have been omitted are of the same order. This is one of the questions this paper will address. Another way to view the effect of wall transpiration is the influence of boundary conditions that vary in the directions parallel to the wall. The first-order asymptotics as developed by Luchini et al. (1991) effectively assumed that the boundary conditions are spatially uniform, and so the wall-normal velocity must be zero at this level of approximation for an impermeable wall at the macroscopic flow scale. However, wall-parallel spatial variations in boundary conditions will allow for effective wall transpiration at the flow scale (Lācis et al., 2020; Wong et al., 2024; Hao & García-Mayoral, 2025). Further, we will also see that order-by-order matching in physical space does not require linearity between the macroscopic forcing and the microscopic response, with no ad-hoc modelling nor tuning parameters.
@@
The exercise recovers the first-order protrusion heights (Bechert & Bartenwerfer, 1989; Luchini et al., 1991) and then extends them to higher orders, which contain nonlinear terms that reflect the nonlinearity of Navier–Stokes equations. The general idea that made this possible is to apply matched inner and outer asymptotic expansions in the classical sense of van Dyke (1964) but then, instead of solving the inner and outer equations in a back-and-forth cascade, represent the outer solution by a formal Cauchy series and only (numerically) solve the inner equations. This allows an equivalent boundary condition to be formulated, at any desired order of accuracy, that completely replaces the inner dynamics by its projection on a flat virtual wall. First we start with the 2D Laplace equation to demonstrate the mechanics of the process, and then apply it with increasing complexity to Stokes and then the 3D Navier–Stokes equations. The main operational results, at a first reading that skips the mathematics involved, are collected in the third-order equivalent boundary condition (7.1) and the table of its coefficients for five typical wall textures, table 1.
@@
Figure 1: Setup: $s^{*}$ is the riblet period/spacing, $L^{*}\gg s^{*}$ is the upper boundary condition plane (measured from the tips at $z^{*}=0$) where flow on length scales O$(L^{*})$ is imposed.
@@
in a 2D strip characterized by a periodic arrangement of protruberances or cavities (riblets) of period $s^{*}$ on one wall (figure 1) and an opposing infinite flat wall at far-away distance $L^{*}\gg s^{*}$. Coordinates are named $y^{*}$ along the wall and $z^{*}$ normal to the wall with an eye to later three-dimensional applications. A superscript is used to mark these quantities as dimensional, as opposed to plain letters used further down to denote dimensionless variables. Also, later this will become a flow problem, but for the time being let the unknown $\phi^{*}$ be any scalar quantity that obeys the Laplace equation, say a temperature field. For definiteness let us also impose a Dirichlet boundary condition ($\phi^{*}=0$) on the $s^{*}$-periodic riblet wall, located near $z^{*}=0$, and a constant Neumann boundary condition ($\phi^{*}_{z^{*}}={}$const.) on the flat wall located at $z^{*}=L^{*}$.
@@ 2.1 Wavenumber domain (constant outer conditions)
where $\beta^{*}_{1}=2\pi/s^{*}$. Far from the periodic wall the mean ($n=0$) component is given in (4) of Luchini et al. (1991),
@@
whereas all the $n\neq 0$ components decay exponentially as $\exp(-\beta^{*}_{1}z^{*})$ or faster. It follows that in most of the domain (precisely for $\beta^{*}_{1}z^{*}\gg 1$, or $z^{*}\gg s^{*}$) the solution has a linear profile as if the wall were flat, and appears to cross zero at a virtual origin $z^{*}=-h^{*}$. This virtual origin, or its distance from an arbitrary reference location which often is taken to be the riblet tip, by definition is the protrusion height, so named by Bechert & Bartenwerfer (1989) who first introduced and calculated it for the Laplace equation applied to longitudinal flow past riblets. Since the governing equation is linear, and the whole solution $\phi^{*}$ is proportional to the strength of the imposed flux $\hat{\phi}^{*}_{0,z^{*}}(L^{*})$, the protrusion height is independent of $\hat{\phi}^{*}_{0,z^{*}}(L^{*})$ and can be obtained once and for all for a given riblet geometry from a single numerical computation, several examples of which were given in Luchini et al. (1991). In addition, given the scale invariance of the Laplace equation, $h^{*}$ is proportional to $s^{*}$ and can be written as $h^{*}=h_{1}s^{*}$, where $h_{1}$ is a dimensionless number calculated for period equal to unity.
@@ 2.2 Physical domain (slowly varying outer conditions)
In order to introduce matched asymptotic expansions, in the style in which they are usually introduced in boundary layers (van Dyke, 1964), we need inner coordinates $Y,Z$, made dimensionless with the riblet period $s^{*}$, and outer coordinates $y=y_{0}+\varepsilon Y,z=\varepsilon Z$ made dimensionless with the outer scale $L^{*}$ in a neighbourhood of a generic boundary point $(y_{0},0)$. The, now dimensionless, dependent variable $\phi$ can be expanded in an inner series
@@
In addition, $\varPhi_{i}$ will obey a Dirichlet boundary condition on a curve $Z=F(Y)$, $\varPhi_{i}[Y,F(Y)]=0$, with the riblet shape $F(Y)$ periodic of period $1$ (because $Y$ has been made dimensionless with the period), and $\phi_{i}$ a variable Neumann boundary condition at the outer wall $z=1$:
@@
In order to match the inner to the outer solution in a neighbourhood of a generic position $y_{0}$, we can set $y=y_{0}+\varepsilon Y$, $z=\varepsilon Z$, and expand each term of the outer series (2.5) in a 2D Taylor series in $\varepsilon Y$ and $\varepsilon Z$:

### v2

@@ Abstract
The description of riblets and other drag-reducing devices has long used the concept of longitudinal and transverse protrusion heights, both as a means to predict the drag reduction itself and as equivalent boundary conditions to simplify numerical simulations by transferring the effect of riblets onto a flat virtual boundary. The limitation of this idea is that it stems from a first-order approximation in the riblet-size parameter $s^{+}$, and as a consequence it cannot predict other than a linear dependence of drag reduction upon $s^{+}$; in other words, the initial slope of the drag-reduction curve. Here the concept is extended to a full asymptotic expansion using matched asymptotics, which consistently provides higher-order protrusion coefficients and higher-order equivalent boundary conditions on a virtual flat surface. While the majority of our results, though nonlinear in $s^{+}$, remain linear in velocity, and therefore we shall not directly address the shape of the drag-reduction curve, this procedure will also allow us to explore the way nonlinearities of the Navier-Stokes equations first enter the $s^{+}$-expansion, with somewhat surprising negative results.
@@ 1 Introduction
Surfaces that are not smooth alter the flow-induced transfer of momentum, heat and mass (Jiménez, 2004; Bottaro, 2019; García-Mayoral et al., 2019; Chung et al., 2021). Examples include riblets (García-Mayoral and Jiménez, 2011a), shark denticles (Lang et al., 2008; Savino and Wu, 2024), seal fur (Itoh et al., 2006; Gómez-de-Segura and García-Mayoral, 2019), superhydrophobic or liquid-infused surfaces (Van Buren and Smits, 2017; García-Cartagena et al., 2019), roughness (Busse et al., 2017; Thakkar et al., 2018; Flack and Schultz, 2010, 2014) and wavy surfaces (Charru et al., 2013; Thorsness et al., 1978). Understanding the relationship between form (shape, topography, morphology) and function (alteration of transfer properties) of these surfaces is not trivial, generally requiring challenging laboratory experiments (Abrams and Hanratty, 1985; Bechert et al., 1997; Schultz and Flack, 2013) or numerical simulations (Busse et al., 2017; García-Mayoral and Jiménez, 2011b). However, in regimes where a small parameter exists, inroads can be made with asymptotic analyses.
@@
As discussed in Luchini (2013), surface texture characterized by a small scale can be analysed in two separate distinguished limits, one of which is a shallow-roughness limit where height of the texture tends to zero while its horizontal scales stay fixed. These are shallow or wavy surfaces (Luchini and Charru, 2019; Abrams and Hanratty, 1985). The other limit is the small-roughness limit where all three dimensions tend to zero proportionally to each other, i.e. the texture stays geometrically self-similar, e.g. riblets (García-Mayoral and Jiménez, 2011b; Endrikat et al., 2021; Modesti et al., 2021) and rough surfaces (Thakkar et al., 2018; Sharma and García-Mayoral, 2020) with small surface features compared to flow features. In this paper we shall be concerned with the small-roughness limit. In particular, we will develop a higher-order matched asymptotic expansion that encapsulates the relationship between flow alteration and surface details via equivalent (homogenised) boundary conditions. Apart from their direct use for informing design (e.g. García-Mayoral and Jiménez, 2011a), it is hoped that such boundary conditions circumvent the need to resolve the small surface features in flow simulations, easing a bottleneck, which then enables routine computations and subsequent analysis towards overall drag savings, for example (Lācis et al., 2020; Ahmed and Bottaro, 2025).
@@
In the context of riblets, Bechert and Bartenwerfer (1989) first developed the idea of the longitudinal protrusion height, which can also be interpreted as a slip length (cf. Bottaro, 2019). When applied in the turbulent-flow regime, it predicts the wall-normal distance below a reference plane (often taken as the plane of the riblet tips) where the mean (longitudinal) flow extrapolates linearly to zero. In this sense, the theory predicts that the mean flow is shifted by riblets, an idea which has been confirmed in numerical simulations (García-Mayoral and Jiménez, 2011b). In the presence of three-dimensional turbulence, it was also necessary to consider the other (non-longitudinal) flow components, and their own shift analogous to the one of the mean flow. This was explored by Luchini et al. (1991), who identified the shift in turbulent fluctuations to be the lateral- or cross-flow (spanwise to the mean) protrusion height. Only after both shifts are identified can the drag change that modifies the inner layer of wall turbulence be explained and extrapolated to full scale (Luchini, 1996). Specifically, the drag change, in the linearized limit, is proportional to the difference between the longitudinal and cross-flow protrusion heights, which is independent of the reference plane from which the protrusion heights are measured.
The protrusion heights (Luchini et al., 1991) describe well the linear asymptotic change in drag for vanishingly small roughness in turbulent flow (Wong et al., 2024), however, the effect of wall transpiration, i.e. wall-normal component of motion, needs to also be accounted for where small roughness features are no longer vanishingly small (Gómez-de-Segura et al., 2018; Lācis et al., 2020; Ibrahim et al., 2021; Sudhakar et al., 2021). Transpiration is akin to a wall-normal shift in “virtual origin” of the wall-normal velocity component (Ibrahim et al., 2021). In cases even beyond the linear asymptotic regime, turbulence is often unmodified, except for a wall-normal shift, itself a nonlinear combination of shifts of both spanwise and wall-normal velocities (Ibrahim et al., 2021). For riblets that are comparable in size to the quasi-streamwise vortices of turbulence (i.e. no longer vanishingly small), a physically based so-called viscous vortex model was developed by Wong et al. (2024) to allow prediction for the shift in turbulence, consistent with non-zero wall-normal (transpiration) component.
Another approach is the transpiration-resistance model of Lācis et al. (2020); Khorasani et al. (2022), which are essentially homogenised boundary conditions for the macroscopic flow, with coefficients obtained from microscopic flow calculations performed in a single repeating unit of surface texture. A similar homogenisation approach is also pursued by Zampogna et al. (2019). As shown by Bottaro and Naqvi (2020) and Sudhakar et al. (2021), the transpiration part appears at the second order of an asymptotic expansion. (The word “order”, here and everywhere in this paper, denoting the power of the expansion parameter $\varepsilon$ with which various quantities approach the limit of $\varepsilon\rightarrow 0$.) Third-order boundary conditions are developed by Bottaro and Naqvi (2020). However, the focus of Bottaro and Naqvi (2020) and Sudhakar et al. (2021) are both on two-dimensional, two-component flow; a three-dimensional, three-component third-order boundary condition has not yet been developed. In terms of range of applicability, a limitation is the absence of advection at lower orders, and so an innovation is a linear Oseen-like approximation (Ahmed and Bottaro, 2024, 2025), a technique already used for porous-medium homogenisation (Zampogna and Bottaro, 2016; Wittkowski et al., 2024). Another approach in this vein is the slip-transpiration-vortex model (Bottaro et al., 2025) that approximates the effect of advection.
@@
Although the effect of wall transpiration has been recognised as a second-order effect and modelled to varying degrees (Lācis et al., 2020; Ibrahim et al., 2021; Wong et al., 2024; Ahmed and Bottaro, 2025), it is unclear at which orders do unsteadiness, nonlinear advection, etc. appear and, conversely, what effects appear at the next (third) order. This is one of the questions this paper will address. We will also see that the effect of shape asymmetry appears to have been missed. Another way to view the effect of wall transpiration is the influence of boundary conditions that vary in the directions parallel to the wall. The first-order asymptotics as developed by Luchini et al. (1991) effectively assumed that the boundary conditions are spatially uniform, and so the wall-normal velocity must be zero at this level of approximation for an impermeable wall at the macroscopic flow scale. However, wall-parallel spatial variations in boundary conditions will allow for effective wall transpiration at the flow scale (Lācis et al., 2020; Wong et al., 2024; Hao and García-Mayoral, 2025). Further, we will also see that order-by-order matching in physical space does not require linearity between the macroscopic forcing and the microscopic response, with no ad-hoc modelling nor tuning parameters.
The approach we adopt here will be a rational asymptotic expansion in powers of a length ratio $\varepsilon$. Any asymptotic expansion critically depends on the quantities that are kept constant during the limiting process of $\varepsilon\rightarrow 0$, i.e. on the scaling of quantities that are not explicitly involved in the process (a bit like partial derivatives depend on what other variables are implicitly kept constant during the differentiation). This is what is meant by “distinguished limit”, and it has to be specified carefully. Flow past a riblet surface can be considered as having three characteristic lengths: the riblet size $s^{*}$ (chosen to be represented by their period), the distance to the opposing wall $L^{*}$ and a dynamic length which will be taken to be the viscous length $\ell^{*}=\nu^{*}/u^{*}_{\tau}$. These three lengths will combine in various ways to determine the distinguished limit, according to their ordering. Another defining feature of the distinguished limit that we adopt here is that all three components of velocity are assumed to scale proportionally to each other, i.e. share the same reference velocity $u^{*}_{R}=u^{*}_{\tau}$. This is ordinarily the case in a near-wall turbulent flow but not, for instance, in a laminar boundary layer.
@@
The exercise recovers the first-order protrusion heights (Bechert and Bartenwerfer, 1989; Luchini et al., 1991) and then extends them to higher orders, which contain nonlinear terms that reflect the nonlinearity of Navier–Stokes equations. The general idea that made this possible is to apply matched inner and outer asymptotic expansions in the classical sense of van Dyke (1964) but then, instead of solving the inner and outer equations in a back-and-forth cascade, represent the outer solution by a formal Cauchy series and only (numerically) solve the inner equations. This allows an equivalent boundary condition to be formulated, at any desired order of accuracy, that completely replaces the inner dynamics by its projection on a flat virtual wall. First we start with the 2D Laplace equation to demonstrate the mechanics of the process, and then apply it with increasing complexity to Stokes and then the 3D Navier–Stokes equations. The process will gradually lead to our main operational result exhibited in §7, the third-order equivalent boundary condition (7.1) and the table of its coefficients for six typical wall textures, table 1.
@@
[…]
@@
in a 2D strip characterized by a periodic arrangement of protuberances or cavities (riblets) of period $s^{*}$ on one wall (figure 1) and an opposing infinite flat wall at far-away distance $L^{*}\gg s^{*}$. Coordinates are named $y^{*}$ along the wall and $z^{*}$ normal to the wall with an eye to forthcoming three-dimensional applications. A superscript is used to mark these quantities as dimensional, as opposed to plain letters used further down to denote dimensionless variables. Just as well, a notation such as $\phi^{*}_{z^{*}}$ denotes a dimensional derivative of $\phi^{*}$ with respect to $z^{*}$. Also, later this will become a flow problem, but for the time being let the unknown $\phi^{*}$ be any scalar quantity that obeys the Laplace equation, say a temperature field. For definiteness let us also impose a Dirichlet boundary condition ($\phi^{*}=0$) on the $s^{*}$-periodic riblet wall, located near $z^{*}=0$, and a constant Neumann boundary condition ($\phi^{*}_{z^{*}}={}$const.) on the flat wall located at $z^{*}=L^{*}$. A Neumann boundary condition at the wall would require nontrivial modifications to what follows, however we shall only need a Dirichlet condition in the subsequent application to fluid flow.
@@ 2.1 Wavenumber domain (constant outer conditions)
where $\beta^{*}_{1}=2\pi/s^{*}$, $\mathrm{i}$ is the imaginary unit, and $n$ numbers the harmonics. Far from the periodic wall the mean ($n=0$) component is given in (4) of Luchini et al. (1991),
@@
whereas all the $n\neq 0$ components decay exponentially as $\exp(-\beta^{*}_{1}z^{*})$ or faster. It follows that in most of the domain (precisely for $\beta^{*}_{1}z^{*}\gg 1$, or $z^{*}\gg s^{*}$) the solution has a linear profile as if the wall were flat, and appears to cross zero at a virtual origin $z^{*}=-h^{*}$. This virtual origin, or its distance from an arbitrary reference location which often is taken to be the riblet tip, by definition is the protrusion height, so named by Bechert and Bartenwerfer (1989) who first introduced and calculated it for the Laplace equation applied to longitudinal flow past riblets. Since the governing equation is linear, and the whole solution $\phi^{*}$ is proportional to the strength of the imposed flux $\hat{\phi}^{*}_{0,z^{*}}(L^{*})$, the protrusion height is independent of $\hat{\phi}^{*}_{0,z^{*}}(L^{*})$ and can be obtained once and for all for a given riblet geometry from a single numerical computation, several examples of which were given in Luchini et al. (1991). In addition, given the scale invariance of the Laplace equation, $h^{*}$ is proportional to $s^{*}$ and can be written as $h^{*}=h_{1}s^{*}$, where $h_{1}$ is a dimensionless number calculated for period equal to unity.
@@ 2.2 Physical domain (slowly varying outer conditions)
In order to introduce matched asymptotic expansions, in the style in which they are usually introduced in boundary layers (van Dyke, 1964), we need inner coordinates $Y,Z$, made dimensionless with the riblet period $s^{*}$, and outer coordinates $y=y_{0}+\varepsilon Y,z=\varepsilon Z$ made dimensionless with the outer scale $L^{*}$ in a neighbourhood of a generic boundary point $(y_{0},0)$. As a mnemonic, big $Z$ is a numerically big dimensionless ordinate, $z^{*}$ divided by the inner reference length, and small $z$ is a small dimensionless ordinate, $z^{*}$ divided by the outer reference length, so that $z=\varepsilon Z$. The, now dimensionless, dependent variable $\phi$ can be expanded in an inner series
@@
In addition, $\varPhi_{i}$ will obey a Dirichlet boundary condition on a curve $Z=F(Y)$, $\varPhi_{i}[Y,F(Y)]=0$, with the riblet shape $F(Y)$ periodic of period $1$ (because $Y$ has been made dimensionless with the period), and $\phi_{i}$ will obey a variable Neumann boundary condition at the outer wall $z=1$:
@@
In order to match the inner to the outer solution in a neighbourhood of a generic boundary position $(y_{0},0)$, we can set $y=y_{0}+\varepsilon Y$, $z=\varepsilon Z$, and expand each term of the outer series (2.5) in a 2D Taylor series in $\varepsilon Y$ and $\varepsilon Z$: